A quadratic equation, in general form, has positive integercoefficients and two real roots. If one of the roots is -2/5 and the difference between the roots is 2, find the equation.

Found 2 solutions by stanbon, josgarithmetic:
Answer by stanbon(75887)

A quadratic equation, in general form, has positive integercoefficients and two real roots. If one of the roots is -2/5 and the difference between the roots is 2, find the equation.
---------
If -2/5 is a root, (5x+2) is a factor.
------
If -2/5+2 = -2/5+10/5 = 8/5 is a root, (5x-8) is a factor.
-----
Equation:
f(x) = (5x+2)(5x-8) = 25x^2-30x-16
-----------------------------------------
Cheers,
Stan H.
-----------------

Answer by josgarithmetic(39617)